OpenStudy (anonymous):

evaluate f(x)=-2x-5 for x=3

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

Okay so you just plug in x=3 in the first equation you stated!

OpenStudy (anonymous):

f(x) = -2 (3) - 5 = -6 -5 = -11

OpenStudy (anonymous):

ok thanks alot

OpenStudy (anonymous):

Yup!